OpenStudy (anonymous):

Find the equation of the tangent to the following parabola 3y^2+16x=0 ; two perpendicular tangents one of the two point of contact being (-3,4)

OpenStudy (anonymous):

Hi!!

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Okay, I will reinterpret the question for you.

OpenStudy (solomonzelman):

A horizontal parabola \(3y^2+16x=0\) will have the reflection of what is above the x-axis, below the x-axis. (Read this sentence carefully) Thus, at x=-3 there are going to be two tangent lines. And you have to choose the tangent line to the curve at (-3,4) and (-3,-4), BUT you have to choose (-3,4).

OpenStudy (solomonzelman):

So our function can be also written as: \(3y^2+16x=0\) \(3y^2=-16x\) \(y^2=-\frac{16}{3}x\) \(y=\color{red}{+} \sqrt{-\frac{16}{3}x}\)

OpenStudy (solomonzelman):

I am choosing only the top half of the functio because the point that we are considering is on the top half of the function.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

So, \(f(x)= \sqrt{-\frac{16}{3}x}\) find the equation of the tangent line, at the point (-3,4).

OpenStudy (solomonzelman):

What would be the first thing to do here?

OpenStudy (anonymous):

find the slope?

OpenStudy (solomonzelman):

yes, and to do that you have to find the derivative.

OpenStudy (solomonzelman):

What is the derivative of our function ?

OpenStudy (anonymous):

y'=1/2*-16/3x^-1/2 y'=(-8/3)x^-1/2 y'=-8/3sq.rtx

OpenStudy (solomonzelman):

correct yourself.

OpenStudy (anonymous):

\[y'=\frac{ -8 }{ 3\sqrt{x} }\]

OpenStudy (solomonzelman):

yes, but your quantity inside the square root is?

OpenStudy (solomonzelman):

\(\color{#226633}{f(x)= \dfrac{-8}{3\sqrt{-\dfrac{16}{3}x}}}\)

OpenStudy (solomonzelman):

Now, the slope at x=-3 is f\('\)(-3). Can you find that ?

OpenStudy (anonymous):

f(-3)=-2/3

OpenStudy (solomonzelman):

yes, f\('\)(-3)=-2/3

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/9yd8w5xuyf